# Leveraging Synthetic Audio Data for End-to-End Low-Resource Speech Translation

### v1

@@ 5 Evaluation and Results
As Table 2 shows, after fine-tuning Whisper Medium on both the authentic and synthetic audio data (Model B), there are consistent improvements across all metrics compared to when we fine-tuned it on the authentic audio data only (Model A). Moreover, Model B++ that uses three synthetic datasets outperforms Model B that uses only two synthetic datasets. This demonstrates that augmented authentic audio data with high-quality synthetic audio data can enhance end-to-end speech translation systems, especially for low-resource language like Irish.
Model C uses the same data as Model B as well as two augmented versions, one version that applies basic VAD, removing low-amplitude samples (cf. Section 3.3.1) and another version that injects white background noise into the data (cf. Section 3.3.2). Although Model C that uses noise and VAD augmented data still outperforms Model A that uses authentic data only, both Model B and B++ that combines authentic data with synthetic data outperform Model C.
@@ A.1 Inference VAD
Table 3: Fast-Whisper VAD default values.
@@ A.2 Training Warm-up Ratio
Table 4: Comparing diverse values of warm-up ratio at training time. Ratios are out of 3000 steps. Hence, 0.01 and 0.03 correspond to 30 steps and 90 steps, respectively. The results shown here are with VAD at inference only, with CTranslate2 default VAD arguments. The highest score in each group is displayed in a bold font.
@@
Table 5: Comparing diverse values of warm-up ratio at training time. Ratios are out of 3000 steps. Hence, 0.01 and 0.03 correspond to 30 steps and 90 steps, respectively. The results shown here are without VAD at either training or inference. The highest score in each group is displayed in a bold font.

### v2

@@ 5 Evaluation and Results
As Table 2 shows, after fine-tuning Whisper Medium on both the authentic and synthetic audio data (Model B), there are consistent improvements across all metrics compared to when we fine-tuned it on the authentic audio data only (Model A). Moreover, Model B++ that uses three synthetic datasets outperforms Model B that uses only two synthetic datasets. This demonstrates that augmented authentic audio data with high-quality synthetic audio data can enhance end-to-end speech translation systems, especially for low-resource languages like Irish.
Model C uses the same training data as Model B as well as two augmented versions, one version that applies basic VAD, removing low-amplitude samples (cf. Section 3.3.1) and another version that injects white background noise into the data (cf. Section 3.3.2). Although Model C that uses noise and VAD augmented data still outperforms Model A that uses authentic training data only, both Model B and B++ that combines authentic data with synthetic data outperform Model C.
@@ A.1 Inference VAD
Table 3: Default VAD values of Faster-Whisper.
@@ A.2 Training Warm-up Ratio
Table 4: Comparing diverse values of warm-up ratio at training time. Ratios are out of 3000 steps. Hence, 0.01 and 0.03 correspond to 30 steps and 90 steps, respectively. The results here are with VAD at inference time, using the default VAD arguments of Faster-Whisper. The highest score in each group is displayed in a bold font.
@@
Table 5: Comparing diverse values of warm-up ratio at training time. Ratios are out of 3000 steps. Hence, 0.01 and 0.03 correspond to 30 steps and 90 steps, respectively. The results here are without VAD at inference time. The highest score in each group is displayed in a bold font.